OpenStudy (anonymous):

Someone Please Help!!! Find the radian measure of an angle of -280 degrees

OpenStudy (anonymous):

Do you have connections academy?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

to find the radian measure multiply -280 * pi radians/180

OpenStudy (anonymous):

-280 = -270 - 10 = -(270 + 10) = -(2(pi)/3 + 10)

OpenStudy (anonymous):

or like the dude said, multiply with (pi)/180

OpenStudy (anonymous):

-14pi/9

OpenStudy (anonymous):

alright. I'm still slightly confused. I have to be shown things to understand better, but i somewhat understand. Can you help me with some other problems?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

find the degree measure of an angle of 3(pi)/5 radians.

OpenStudy (anonymous):

Thanks you so much!:)

OpenStudy (anonymous):

you multiply 3pi/5radians* 180/pi radians to get 108

OpenStudy (anonymous):

okay, i definitely understand that one. Find the exact values of cos(3(pi)/4 radians) and sin (3(pi)/4 radians).

OpenStudy (anonymous):

I still don't understand that one..but its -sqrt2/2, sqrt2/2

OpenStudy (anonymous):

Okay, i will just have to ask my teacher about that one later. Use the given circle. Find the length s to the nearest tenth.

OpenStudy (anonymous):

length of intercepted arc = radius*central angle

OpenStudy (anonymous):

so 5 * 2pi/3 = 10.5

OpenStudy (anonymous):

Understandable. When you explain this and i understand it, it makes me feel slightly dumb for not understanding when i was reading it in the book. lol Last one. A weather satellite in circular orbit around Earth completes one orbit every 5 hours. The radius of Earth is about 6,400 km and the satellite is positioned 4,700 km above the Earth. How far does the satellite travel in 1 hour? Round your answer to the nearest kilometer.

OpenStudy (anonymous):

Thats why i take notes lol even if i don't understand it, it is written in front of me I don't understand how to do this last one but the answer is 13,949

OpenStudy (anonymous):

lol yeah, it is just gibberish to me either way.. lol Thanks so much!!:)

OpenStudy (anonymous):

B C C A C 100%